OpenStudy (awy):

Assume that a sample is used to estimate a population proportion mu or mean. Find the margin of error M.E. that corresponds to a sample of size 18 with a mean of 20.1 and a standard deviation of 13.5 at a confidence level of 90%.

OpenStudy (awy):

so i guess for this problem i have to use z(alpha/2) * (standard dev/sqrt(n)) right?

OpenStudy (awy):

because i'm tryig to find my margin of error

OpenStudy (kirbykirby):

A confidence interval for the mean (with a small sample) is calculated as: \[\bar{x} \pm t * \frac{s}{\sqrt{n}}\] but can also be written as: \[ \bar{x} \pm \text{margin of error}\] so the margin of error is \[t* \frac{s}{\sqrt{n}} \] where \(t\) is the critical value for a 90% confidence interval from the t-distribution with degrees of freedom = \(n-1\), and \(s\) is the standard deviation and \(n\) is the sample size.

OpenStudy (awy):

so, its 1.645 * (13.5/sqrt(18)) = 5.234

OpenStudy (kirbykirby):

You can use a normal distribution, with a critical \(Z\) values instead of a critical \(t\) value if the standard deviation given is a known population standard deviation (rather than a sample standard deviation, but it's not "clear" from the question, and usually the pop. std. dev. is not given). Otherwise, the sample size is small, so usually small is considered n = 30 (but this might depend on your prof), in which case the t-distribution will be better

OpenStudy (kirbykirby):

small is considered n <= 30*

OpenStudy (awy):

oh, in my textbook it says N<30

OpenStudy (awy):

n*

OpenStudy (kirbykirby):

ok, then yes I suggest you use a t-table for the critical value (rather than using 1.645 which is form a normal distribution table)

OpenStudy (awy):

so, 1.740

OpenStudy (awy):

1.740 * (13.5/sqrt(18)) = 5.5

OpenStudy (awy):

so anytime n is less than or equal to 30, i have to use the t table?

OpenStudy (kirbykirby):

Yes, UNLESS the problem specifically tells you that they gave you a -population- standard deviation (rather than a -sample- standard deviation).

OpenStudy (awy):

so you mean like the pop. is not normally distributed and n < or = to 30?

OpenStudy (awy):

ty for the help.

OpenStudy (awy):

Also another question, how would i know in the problem when to use the t distributions and the standard normal?

OpenStudy (kirbykirby):

1) If the population standard deviation is given, and the sample size is large (n >= 30), use a normal distribution 2) If the population standard deviation is given, and the sample size is small (n < 30), use the normal distribution 3) If the population standard deviation is NOT given (the sample std. deviation is given instead), and the sample size is large (n >= 30), use the normal distribution (however, using the t distribution is actually more accurate to use), but the normal approximation is usually okay to use. 4) If the population standard deviation is NOT given (the sample std. deviation is given instead), and the sample size is small (n < 30), use the t distribution for degrees of freedom n-1 ... note that cases 3 and 4 are usually what occur... it is very rare and unlikely that you will have population standard deviations at your disposal like in cases 1 and 2. But it's something to be aware of in the problem.

OpenStudy (awy):

Awesome! thank you very much for the help and info. It helped a lot.

OpenStudy (kirbykirby):

your welcome! :)

OpenStudy (kirbykirby):

Maybe an easier summary for the conditions is: